what is the weight of a 236 kg object on the surface of venus if the radius of venus is 6,000 km and the mass of venus is a 4.87x10^24 kg?
1
Expert's answer
2012-11-05T10:07:08-0500
weight is equal to gravity force, which is
F=G*m1*m2/r^2
F=6.67*10^{-11}*236*4.87*10^{24}/(6000000^2)=212.9 kg